The length of a rectangle is 3cm longer than its width. If the perimeter of the rectangle is 26cm , find its area.
Answer by mananth(16946)

The length of a rectangle is 3cm longer than its width. If the perimeter of the rectangle is 26cm , find its area.
Let the width be x
the length will be x+3
perimeter of rectangle = 2L+2W
= 2*x+2*(x+3)
2x+2x+6 =26
4x=26-6
x=20/4
x=5
Length will be x+3 = 5+3 =8
Area of rectangle = L * W
= 5*8 = 40 square cms.